# Combating Confirmation Bias: A Unified Pseudo-Labeling Framework for Entity Alignment

### v2

@@ 3.2.2 Pseudo-Label Ensembling
Formally, given $m$ sets of pseudo-labeled alignments $\{\widehat{\mathcal{S}}_{t}|t=1,...,m\}$ inferred from every $m$ consecutive iterations, we generate pseudo-labeled alignments that are commonly selected in all of $\{\widehat{\mathcal{S}}_{t}|t=1,...,m\}$:
$\widehat{\mathcal{S}}=\cap_{t=1}^{m}\widehat{\mathcal{S}}_{t}=\bigl{\{}(e_{i},e_{j})|\sum_{t=1}^{m}\mathbbm{1}\bigl{(}(e_{i},e_{j})\in\widehat{\mathcal{S}}_{t}\bigl{)}=m,e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}\bigl{\}},$ (16)
@@
where $\mathbbm{1}(\cdot)$ is a binary indicator function. To decorrelate the dependency among $m$ consecutive iterations, model parameters are randomly reinitialized during each of these iterations [5].
By focusing on consensus alignments across iterations, our self-ensembling approach is expected to have higher pseudo-labeling precision than using the predictions from the most recent iteration alone. This approach is related to self-ensembling techniques like temporal ensembling [14], which uses exponential moving average smoothing to aggregate model predictions across iterations and adds regularization to the loss function to improve prediction consistency on unlabeled data. Our approach shares the same objective but provides a simple yet effective solution, which also supports UPL-EA’s utility to integrate with other EA models (as empirically demonstrated in Section 4.3.2 and  4.3.3).
@@
Finally, the refined pseudo-labeled alignments $\widehat{\mathcal{S}}$ are used to augment prior seed alignments $\mathcal{S}_{0}$ as follows:
${\color[rgb]{0,0,0}\mathcal{S}_{0}\leftarrow\mathcal{S}_{0}\cup\widehat{\mathcal{S}}}.$ (17)
@@
The augmented set $\mathcal{S}_{0}$ includes a considerable number of reliable pseudo-labeled alignments, which in turn strengthens subsequent model training.
3.3 Overall Workflow
The training process of UPL-EA is summarized in Algorithm 2. In Steps 2-5, the EA model and OT-Based pseudo-labeling alternately reinforce each other to learn informative entity embeddings. In Steps 6-7, the pseudo-label ensembling is performed every $m$ iterations to augment prior seed alignments with reliably selected pseudo-labeled alignments for subsequent training. Finally, the learned entity embeddings are returned as the output. For alignment inference, the learned entity embeddings are used to infer new aligned entity pairs via Algorithm 1.
Input: Two KGs $\mathcal{G}_{1}=\{\mathcal{E}_{1},\mathcal{R}_{1},\mathcal{T}_{1}\}$, $\mathcal{G}_{2}=\{\mathcal{E}_{2},\mathcal{R}_{2},\mathcal{T}_{2}\}$, prior seed alignments $\mathcal{S}_{0}$.
@@ 3.3 Overall Workflow
2       for t in 1 …$m$ do
3             Infer pseudo-labeled alignments $\widehat{\mathcal{S}}_{t}$ via Algorithm 1;
4            Augment seed alignments: $\mathcal{S}\leftarrow\mathcal{S}_{0}\cup\widehat{\mathcal{S}}_{t}$;
5             Learn entity embeddings by minimizing the loss function in Eq. (15);
6      Ensemble pseudo-labeled alignments $\widehat{\mathcal{S}}=\cap_{t=1}^{m}\widehat{\mathcal{S}}_{t}$;
7       Augment prior seed alignments: $\mathcal{S}_{0}\leftarrow\mathcal{S}_{0}\cup{\color[rgb]{0,0,0}\widehat{\mathcal{S}}}$;
@@ 6      Ensemble pseudo-labeled alignments $\widehat{\mathcal{S}}=\cap_{t=1}^{m}\widehat{\mathcal{S}}_{t}$;
In Algorithm 2, the time complexity of OT-based pseudo-labeling in Step 3 is $O(|\mathcal{E}^{U}_{1}|\cdot|\mathcal{E}^{U}_{2}|\cdot d)$, where $d$ is the entity embedding dimension, and the time complexity of learning entity embeddings in Step 5 is $O((|\mathcal{E}_{1}\cup\mathcal{E}_{2}|)\cdot d^{2})$. Thus, the time complexity of Algorithm 2 is $O(I\cdot(|\mathcal{E}^{U}_{1}|\cdot|\mathcal{E}^{U}_{2}|\cdot d+(|\mathcal{E}_{1}\cup\mathcal{E}_{2}|)\cdot d^{2}))$, where $I$ is the maximum number of iterations.
@@ 4.1.1 Datasets
Cross-lingual Datasets. DBP15K [38] is a widely used benchmark dataset for cross-lingual entity alignment [10, 19, 52, 62]. It includes three cross-lingual KG pairs extracted from DBpedia, each containing two KGs built upon English and another language (Chinese, Japanese, or French), with 15,000 aligned entity pairs per dataset. SRPRS [11] is a more recent benchmark dataset characterized by sparser connections [11] that are extracted from DBpedia. SRPRS comprises two cross-lingual KG pairs, each with two KGs in English and French/German, and also includes 15,000 aligned entity pairs.
@@ 4.1.2 Experimental Setup
For fair comparisons, we follow the conventional 30%-70% split ratio to randomly partition training and test data on all datasets. We use semantic meanings of entity names to construct entity features. On DBP15K with relatively larger lingustic barriers, we first use Google Translate to translate non-English entity names into English, then look up 768-dimensional word embeddings pre-trained by BERT [9] with English entity names to form entity features. On SRPRS and DPB-YG-15K, we directly look up word embeddings without translation. As each entity name comprises one or multiple words, we further use TF-IDF to measure the contribution of each word towards entity name representation. Finally, we aggregate TF-IDF-weighted word embeddings for each entity to form its entity feature vector.
The settings of UPL-EA are specified as follows: $K=125$, $\beta=0.5$, $\gamma=1$, $\lambda=10$, and $m=3$. The embedding dimension $d$ is set to 300. For BERT pre-trained word embeddings, we use a PCA-based technique [30] to reduce feature dimension from 768 to 300 with minimal information loss. The number of pseudo-labeling iterations is set to 10, where each iteration contains 10 training epochs for the EA model. We implement our model in PyTorch, using the Adam optimizer with a learning rate of 0.001 on DBP15K and DBP-YG-15K, and 0.00025 on SRPRS. The batch size is set to 256. All experiments are run on a computer with an Intel(R) Core(TM) i9-13900KF CPU @ 3.00GHz and an NVIDIA Geforce RTX 4090 (24GB memory) GPU.
@@ 4.2.1 Results on Cross-lingual Datasets
Baselines and Metrics. We compare UPL-EA with 12 state-of-the-art EA models on cross-lingual datasets: DBP15K and SRPRS, which broadly fall into two categories:
The Results. Table 2 and Table 3 report performance comparisons on DBP15K and SRPRS, respectively. This set of results are reported with 30% prior seed alignments used for training. The asterisk (*) indicates that semantic meanngs of entity names are used to construct entity features.
@@
Our results show that UPL-EA significantly outperforms most existing EA models on five cross-lingual KG pairs. In particular, on DBP15KZH_EN, UPL-EA outperforms the second and third performers, CPL-OT and RNM, by over 2% and 10%, respectively, in terms of Hit@1. This affirms the necessity of systematically combating confirmation bias for pseudo-labeling-based entity alignment. It is worth noting that disparities in overall performance can be observed among the five cross-lingual KG pairs, where the lowest accuracy is achieved on DBP15KZH_EN due to its large linguistic barriers. Nevertheless, for the most challenging EA task on DBP15KZH_EN, UPL-EA yields strong performance gains over other baselines.
@@ 4.2.2 Results on Cross-Source Monolingual Datasets
Baselines and Metrics. On monolingual dataset DBP-YG-15K, we compare UPL-EA with five EA models, including BootEA [39], RDGCN [52], BERT-INT [43], TransEdge [40], and PARIS+ [16]. The results of these baselines are obtained from [16]. BootEA, TransEdge, and UPL-EA use relational triplets only, the same as our setting. PARIS+, RDGCN, and BERT-INT use both relational and attribute triplets. For UPL-EA, we report the average results over five runs.
@@
The Results. Table 4 reports performance comparisons on DBP-YG-15K (OpenEA) and DBP-YG-15K (RealEA) with 30% prior seed alignments used for training. The asterisk “*” indicates that semantic meanings of entity names are used in EA modeling. Our results show that UPL-EA outperforms all five baselines across two KG pairs of DBP-YG-15K. On the OpenEA KG pair, UPL-EA achieves nearly perfect performance with all three metrics to be approximately 1, outperforming the second best baseline by more than 2% on $F_{1}$ score. On the RealEA KG pair of DBP-YG-15K, even with duplicated entities in each KG [16], UPL-EA performs competitively with an over 3% improvement on $F_{1}$ score compared to the second best baseline.
@@ 4.3.1 Effectiveness of Different Components
To assess the importance of various components of the proposed UPL-EA framework, we conduct a series of ablation studies on five cross-lingual KG pairs from DBP15K and SRPRS. To provide deeper insights, we undertake ablation studies under two settings: the conventional setting using 30% prior seed alignments and the setting with no prior seed alignments provided. The full UPL-EA model is compared with its ablated variants, with the best performance highlighted by bold. From Table 5 and Table 6, we can see that the full UPL-EA model performs the best in all cases.
@@
Full Model 94.9 97.4 0.96 97.0 98.8 0.98 99.5 99.7 1.00
w.o. Global-lev. Rel. Aggr. 92.6 96.4 0.94 96.4 98.5 0.97 99.3 99.7 1.00
@@
w.o. OT. Pseudo-Labeling 80.1 90.4 0.84 87.6 94.8 0.90 94.4 97.4 0.96
w.o. Pseudo-Label Ensemb. 84.9 90.6 0.87 91.1 95.5 0.93 98.1 99.0 0.98
w.o. OT. & Ensemb. 74.8 87.3 0.80 83.4 93.6 0.87 93.1 97.5 0.95
@@
Full Model 93.1 96.2 0.94 95.7 98.3 0.97 99.2 99.5 0.99
w.o. Global-lev. Rel. Aggr. 91.4 95.3 0.93 95.3 98.1 0.96 99.1 99.7 0.99
@@
w.o. OT. Pseudo-Labeling 66.9 75.5 0.70 76.9 85.2 0.80 91.9 95.2 0.93
w.o. Pseudo-Label Ensemb. 83.0 88.7 0.85 90.1 94.6 0.92 97.8 98.8 0.98
w.o. OT. & Ensemb. 67.1 76.8 0.71 77.1 86.2 0.81 91.4 95.9 0.93
@@
w.o. Global-lev. Rel. Aggr. 97.5 98.9 0.98 97.8 99.2 0.98
@@
w.o. OT. Pseudo-Labeling 93.9 96.6 0.95 94.2 97.5 0.95
w.o. Pseudo-Label Ensemb. 94.6 97.4 0.96 94.8 98.1 0.96
w.o. OT. & Ensemb. 92.7 96.5 0.94 93.6 97.4 0.95
@@
Full Model 97.9 99.2 0.98 97.7 99.2 0.98
w.o. Global-lev. Rel. Aggr. 97.1 98.6 0.98 97.4 98.7 0.98
@@
w.o. OT. Pseudo-Labeling 89.8 93.0 0.91 91.4 95.2 0.93
w.o. Pseudo-Label Ensemb. 94.2 97.0 0.95 94.8 97.7 0.96
w.o. OT. & Ensemb. 89.7 93.1 0.91 91.1 94.9 0.93
  • w.o. Global-lev. Rel. Aggr.: We compare the full UPL-EA model against the variant without using global-level relation aggregation for entity embedding learning. This ablation leads to degraded performance on both settings primarily due to the increase in conflicted misalignments caused by feature over-smoothing.
@@
  • w.o. OT. Pseudo-Labeling: To study the efficacy of OT-based Pseudo-Labeling, we ablate it from the full UPL-EA model. As OT modeling can effectively eliminate a considerable number of Type I pseudo-labeling errors to ensure one-to-one correspondences at each iteration, this ablation results in a profound performance drop across all datasets on both settings.
  • w.o. Pseudo-Label Ensemb.: The ablation of pseudo-label ensembling from UPL-EA also significantly degrades alignment performance, with substantial performance declines observed in both settings, especially on DBP15KZH_EN and DBP15KJA_EN with large linguistic barriers. In contrast, performance drops are less pronounced on DBP15KFR_EN, SRPRSEN_FR, and SRPRSEN_DE with relatively small linguistic barriers. This is attributed to the fact that larger linguistic barriers tend to incur more one-to-one misalignments (Type II pseudo-labeling errors). Our findings confirm that pseudo-label ensembling is crucial for UPL-EA to achieve its full potential, especially when model predictions are less accurate during early training stages.
  • w.o. OT. & Ensemb.: We also analyze the overall effect of ablating both OT modeling and pseudo-label ensembling from the full model. This ablation, conceptually identical to the naive pseudo-labeling strategy [40], has a substantial adverse impact, leading to a dramatic performance drop in all cases. Our results highlight the effectiveness of our proposed UPL-EA framework in systematically combating confirmation bias for pseudo-labeling-based entity alignment.
Note that under the setting with no prior seed alignments, the variant without OT-based pseudo-labeling (w.o. OT. Pseudo-Labeling) has similar performance as compared to the variant completely ignoring confirmation bias (w.o. OT. & Ensemb.). In particular, on DBP15KZH_EN and DBP15KJA_EN, the former variant even performs slightly worse. This is because under the challenging case where there are no prior seed alignments, ablating OT-based pseudo-labeling might incur considerably more conflicted misalignments (Type I pseudo-labeling errors). As a result, it becomes ineffective to refine erroneous pseudo-labeled alignments via ensembling.

### v3

@@ 3.2.2 Parallel Pseudo-Label Ensembling
Formally, in pseudo-labeling iteration $t$, given $M$ sets of pseudo-labeled alignments $\{\widehat{\mathcal{S}}_{t}^{(1)},\widehat{\mathcal{S}}_{t}^{(2)},...,\widehat{\mathcal{S}}_{t}^{(M)}\}$ inferred from $M$ OT-based models independently trained in parallel, we generate the ensembled pseudo-labeled alignments by taking those that are consistently selected across all sets in $\{\widehat{\mathcal{S}}_{t}^{(1)},\widehat{\mathcal{S}}_{t}^{(2)},...,\widehat{\mathcal{S}}_{t}^{(M)}\}$:
${\color[rgb]{0,0,0}\mathcal{S}_{t}=\cap_{m=1}^{M}\widehat{\mathcal{S}}_{t}^{(m)}=\bigl{\{}(e_{i},e_{j})|\sum_{m=1}^{M}\mathbbm{1}\bigl{(}(e_{i},e_{j})\in\widehat{\mathcal{S}}_{t}^{(m)}\bigl{)}=M,e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}\bigl{\}},}$ (11)
@@
where $\mathbbm{1}(\cdot)$ is a binary indicator function. To decorrelate the dependency among $M$ EA models, their model parameters are initialized independently.
By focusing on consensus alignments from multiple OT-based models, our parallel ensembling approach is expected to achieve higher pseudo-labeling precision compared to relying on a single model’s predictions. This strategy relates to consistency-based techniques such as Dual Student [14], which combines loosely coupled predictions from two independently trained models and introduces a stabilization constraint in the loss function to enhance prediction consistency on unlabeled data. While sharing a similar objective, our method offers a simpler yet effective alternative, as empirically demonstrated in Section 4.3.2.
@@
Finally, the ensembled pseudo-labeled alignments $\mathcal{S}_{t}$ are used to augment seed alignments $\mathcal{S}$ as follows:
${\color[rgb]{0,0,0}\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}_{t}}.$ (12)
@@
The augmented seed alignments $\mathcal{S}$ include a considerable number of reliable pseudo-labeled alignments, which in turn strengthen subsequent model training.
Input: Two KGs $\mathcal{G}_{1}=\{\mathcal{E}_{1},\mathcal{R}_{1},\mathcal{T}_{1}\}$, $\mathcal{G}_{2}=\{\mathcal{E}_{2},\mathcal{R}_{2},\mathcal{T}_{2}\}$, prior seed alignments $\mathcal{S}$.
@@
[…]
3             Learn entity embeddings based on $\mathcal{S}$ by minimizing Eq. (2);
4             Infer pseudo-labeled alignments $\widehat{\mathcal{S}}_{t}^{(m)}$ via Algorithm 1;
5
6      Ensemble pseudo-labeled alignments $\mathcal{S}_{t}=\cap_{m=1}^{M}\widehat{\mathcal{S}}_{t}^{(m)}$;
7       Augment seed alignments: $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}_{t}$;
@@
3.3 Overall Training Procedure
The UPL-EA training procedure is summarized in Algorithm 2. In Steps 2-4, $M$ EA models are trained in parallel using seed alignments to obtain informative entity embeddings, which are then passed on to OT-based pseudo-labeling to infer conflict-free pseudo-labeled alignments. In Steps 5-6, parallel pseudo-label ensembling is performed over $M$ models to generate ensembled pseudo-labeled alignments, which are thereafter used to augment seed alignments for subsequent model training. Finally, the learned entity embeddings are returned as the output. For alignment inference, the learned entity embeddings are used to infer new aligned entity pairs via Algorithm 1.
In Algorithm 2, the time complexity of learning entity embeddings in Step 3 is $O((|\mathcal{E}_{1}\cup\mathcal{E}_{2}|)\cdot d^{2})$, where $d$ is the entity embedding dimension, and the time complexity of OT-based pseudo-labeling in Step 4 is $O(|\mathcal{E}^{U}_{1}|\cdot|\mathcal{E}^{U}_{2}|\cdot d)$. Thus, the time complexity of Algorithm 2 is $O(I\cdot(|\mathcal{E}^{U}_{1}|\cdot|\mathcal{E}^{U}_{2}|\cdot d+(|\mathcal{E}_{1}\cup\mathcal{E}_{2}|)\cdot d^{2}))$, where $I$ is the maximum number of pseudo-labeling iterations.
@@ 4.1 Experimental Settings
This section presents the experimental settings, including benchmark datasets used and detailed experimental setup.
@@ 4.1.1 Datasets
Cross-Lingual Datasets. DBP15K [39] is a widely used benchmark dataset for cross-lingual entity alignment [10, 20, 53, 63]. It includes three cross-lingual KG pairs extracted from DBpedia, each containing two KGs built upon English and another language (Chinese, Japanese, or French), with 15,000 aligned entity pairs per dataset. SRPRS [11] is a more recent benchmark dataset characterized by sparser connections [11] that are extracted from DBpedia. SRPRS comprises two cross-lingual KG pairs, each with two KGs in English and French/German, and it also includes 15,000 aligned entity pairs.
@@ 4.1.2 Experimental Setup
For fair comparisons, we follow the conventional 30%-70% split ratio to randomly partition training and test data on all datasets. We use semantic meanings of entity names to construct entity features. On DBP15K with relatively larger linguistic barriers, we first use Google Translate to translate non-English entity names into English, then look up 768-dimensional word embeddings pre-trained by BERT [9] with English entity names to form entity features. On SRPRS and DPB-YG-15K, we directly look up word embeddings without translation. As each entity name comprises one or multiple words, we further use TF-IDF to measure the contribution of each word towards entity name representation. Finally, we aggregate TF-IDF-weighted word embeddings for each entity to form its entity feature vector.
The settings of UPL-EA are specified as follows: $K=125$, $\beta=0.5$, $\gamma=1$, $\lambda=10$, and $M=3$. The embedding dimension $d$ is set to 300. For BERT pre-trained word embeddings, we use a PCA-based technique [31] to reduce feature dimension from 768 to 300 with minimal information loss. The number of pseudo-labeling iterations is set to 9, where each iteration contains 10 training epochs for the EA model. We implement our model in PyTorch, using the Adam optimizer with a learning rate of 0.001 on DBP15K and DBP-YG-15K, and 0.00025 on SRPRS. The batch size is set to 256. All experiments are run on a computer with an Intel(R) Core(TM) i9-13900KF CPU @ 3.00GHz and an NVIDIA Geforce RTX 4090 (24GB memory) GPU.
@@ 4.2.1 Results on Cross-Lingual Datasets
Baselines and Metrics. We compare UPL-EA against 12 state-of-the-art EA models on cross-lingual datasets: DBP15K and SRPRS, which broadly fall into two categories:
The Results. Table 2 and Table 3 report performance comparisons on DBP15K and SRPRS, respectively. This set of results is reported with 30% prior seed alignments used for training. The asterisk (*) indicates that semantic meanings of entity names are used to construct entity features.
Our results show that UPL-EA significantly outperforms most existing EA models on five cross-lingual KG pairs. In particular, on DBP15KZH_EN, UPL-EA outperforms the second and third performers, CPL-OT and RNM, by 2% and over 10%, respectively, in terms of Hit@1. This affirms the necessity of systematically combating confirmation bias for pseudo-labeling-based entity alignment. It is worth noting that disparities in overall performance can be observed among the five cross-lingual KG pairs, where the lowest accuracy is achieved on DBP15KZH_EN due to its large linguistic barriers. Nevertheless, for the most challenging EA task on DBP15KZH_EN, UPL-EA yields strong performance gains over other baselines.
@@ 4.2.2 Results on Cross-Source Monolingual Datasets
Baselines and Metrics. On monolingual dataset DBP-YG-15K, we compare UPL-EA against five EA models, including BootEA [40], RDGCN [53], BERT-INT [44], TransEdge [41], and PARIS+ [17]. The results of these baselines are obtained from [17]. BootEA, TransEdge, and UPL-EA use relational triplets only, the same as our setting. PARIS+, RDGCN, and BERT-INT use both relational and attribute triplets. For UPL-EA, we report the average results over five runs.
@@
The Results. Table 4 reports performance comparisons on DBP-YG-15K (OpenEA) and DBP-YG-15K (RealEA) with 30% prior seed alignments used for training. The asterisk “*” indicates that semantic meanings of entity names are used in EA modeling. Our results show that UPL-EA outperforms all five baselines across two KG pairs of DBP-YG-15K. On the OpenEA KG pair, UPL-EA achieves nearly perfect performance with all three metrics to be approximately 1, outperforming the second best baseline by more than 2% on $F_{1}$ score. On the RealEA KG pair of DBP-YG-15K, even with duplicated entities in each KG [17], UPL-EA performs competitively with an over 5% improvement on $F_{1}$ score compared to the second best baseline.
@@ 4.3 Ablation Studies and Analyses
This section presents a series of ablation studies and in-depth analyses to validate the effectiveness of our proposed UPL-EA framework.
@@ 4.3.1 Effectiveness of Different Components
To assess the importance of various components of the proposed UPL-EA framework, we first conduct a thorough ablation study on five cross-lingual KG pairs from DBP15K and SRPRS. To provide deeper insights, we undertake ablation studies under two settings: the conventional setting using 30% prior seed alignments and the setting with no prior seed alignments provided. The full UPL-EA model is compared with its ablated variants, with the best performance highlighted by bold. From Table 5 and Table 6, we can see that the full UPL-EA model performs the best in all cases.
@@
Full Model 94.7 97.5 0.96 97.4 98.9 0.98 99.4 99.7 1.00
w.o. OT. Pseudo-Labeling 80.1 90.4 0.84 87.6 94.8 0.90 94.4 97.4 0.96
w.o. Parallel Ensembling 84.9 90.6 0.87 91.1 95.5 0.93 98.1 99.0 0.98
w.o. OT. & Ensembling 74.8 87.3 0.80 83.4 93.6 0.87 93.1 97.5 0.95
@@
Full Model 93.0 96.2 0.94 96.0 98.3 0.97 99.2 99.5 0.99
w.o. OT. Pseudo-Labeling 66.9 75.5 0.70 76.9 85.2 0.80 91.9 95.2 0.93
w.o. Parallel Ensembling 83.0 88.7 0.85 90.1 94.6 0.92 97.8 98.8 0.98
w.o. OT. & Ensembling 67.1 76.8 0.71 77.1 86.2 0.81 91.4 95.9 0.93
@@
w.o. OT. Pseudo-Labeling 93.9 96.6 0.95 94.2 97.5 0.95
w.o. Parallel Ensembling 94.6 97.4 0.96 94.8 98.1 0.96
w.o. OT. & Ensembling 92.7 96.5 0.94 93.6 97.4 0.95
@@
Full Model 97.9 99.2 0.98 97.4 99.2 0.98
w.o. OT. Pseudo-Labeling 89.8 93.0 0.91 91.4 95.2 0.93
w.o. Parallel Ensembling 94.2 97.0 0.95 94.8 97.7 0.96
w.o. OT. & Ensembling 89.7 93.1 0.91 91.1 94.9 0.93
@@
  • w.o. OT. Pseudo-Labeling: To study the efficacy of OT-based pseudo-labeling, we ablate it from the full UPL-EA model. As OT modeling can effectively eliminate a considerable number of conflicted misalignments to ensure one-to-one correspondences in pseudo-labeled alignments, this ablation results in a profound performance drop across all datasets on both settings.
  • w.o. Parallel Ensembling: The ablation of parallel pseudo-label ensembling from UPL-EA also significantly degrades alignment performance, with substantial performance declines observed in both settings, especially on DBP15KZH_EN and DBP15KJA_EN with large linguistic barriers. In contrast, performance drops are less pronounced on DBP15KFR_EN, SRPRSEN_FR, and SRPRSEN_DE with relatively small linguistic barriers. This is attributed to the fact that larger linguistic barriers tend to incur more one-to-one misalignments. Our findings confirm that parallel pseudo-label ensembling is crucial for UPL-EA to achieve its full potential, especially when model predictions are less accurate during early training stages.
  • w.o. OT. & Ensembling: We also analyze the overall effect of ablating both OT modeling and parallel pseudo-label ensembling from the full model. This ablation, conceptually identical to the naive pseudo-labeling strategy [41], has a substantial adverse impact, leading to a dramatic performance drop in all cases. Our results highlight the effectiveness of our proposed UPL-EA framework in systematically combating confirmation bias for pseudo-labeling-based entity alignment.
@@
Note that under the setting with no prior seed alignments, the variant without OT-based pseudo-labeling (w.o. OT. Pseudo-Labeling) has similar performance as compared to the variant completely ignoring confirmation bias (w.o. OT. & Ensemb.). In particular, on DBP15KZH_EN and DBP15KJA_EN, the former variant even performs slightly worse. This is because under the challenging case where there are no prior seed alignments, ablating OT-based pseudo-labeling might incur considerably more conflicted misalignments. As a result, it becomes ineffective to filter out erroneous pseudo-labeled alignments via ensembling.
4.3.2 Comparisons with Other Ensembling Methods
[…]
Table 7: Performance of UPL-EA using different ensembling methods.
DBP15KZH_EN DBP15KJA_EN DBP15KFR_EN
Hit@1 Hit@10 MRR Hit@1 Hit@10 MRR Hit@1 Hit@10 MRR
[…]